OpenStudy (anonymous):

Help? Complete the equation. Write your answer in decimal form. 2/4% x 25

jimthompson5910 (jim_thompson5910):

not sure what you want exactly you want 2 divided by 4% or (2/4)% ?

OpenStudy (anonymous):

(2/4)% times 25

jimthompson5910 (jim_thompson5910):

ok 2/4 = 1/2 = 0.5

jimthompson5910 (jim_thompson5910):

so (2/4)% = 0.5% = 0.005

jimthompson5910 (jim_thompson5910):

then you multiply that with 25 to get your answer

OpenStudy (anonymous):

0.125?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

Massive thank you!

jimthompson5910 (jim_thompson5910):

you're welcome